### v1

# Coherent Load Profile Synthesis with Conditional Diffusion for LV Distribution Network Scenario Generation

Abstract
Limited visibility of power distribution network power flows at the low voltage level presents challenges to both distribution network operators from a planning perspective and distribution system operators from a congestion management perspective. Forestalling these challenges through scenario analysis is confounded by the lack of realistic and coherent load data across representative distribution feeders. Load profiling approaches often rely on summarising demand through typical profiles, which oversimplifies the complexity of substation-level operations and limits their applicability in specific power system studies. Sampling methods, and more recently generative models, have attempted to address this through synthesising representative loads from historical exemplars; however, while these approaches can approximate load shapes to a convincing degree of fidelity, the co-behaviour between substations, which ultimately impacts higher voltage level network operation, is often overlooked. This limitation will become even more pronounced with the increasing integration of low-carbon technologies, as estimates of base loads fail to capture load diversity. To address this gap, a Conditional Diffusion model for synthesising daily active and reactive power profiles at the low voltage distribution substation level is proposed. The evaluation of fidelity is demonstrated through conventional metrics capturing temporal and statistical realism, as well as power flow modelling. The results show synthesised load profiles are plausible both independently and as a cohort in a wider power systems context. The Conditional Diffusion model is benchmarked against both naive and state-of-the-art models to demonstrate its effectiveness in producing realistic scenarios on which to base sub-regional power distribution network planning and operations.
keywords:
Load Modelling, Power Systems Modelling, Neural Network Applications, Generative Modelling.
journal: Sustainable Energy, Grids and Networks
[…]
[1]organization=National Manufacturing Institute Scotland, city=Renfrew, postcode=PA3 2EF, country=Scotland
[…]
[2]organization=Department of Electronic and Electrical Engineering, University of Strathclyde, city=Glasgow, postcode=G1 1XW, country=Scotland
[…]
[3]organization=Scottish and Southern Energy Networks, city=Perth, postcode=PH1 3AF, country=Scotland
1 Introduction
Establishing the extent to which low-voltage (LV) distribution feeders are challenged is a key obstacle to achieving decarbonisation of energy end use in the form of heat and transportation. In LV networks, which in Great Britain (GB) is 415V for domestic and light commercial customers [1], visibility has become increasingly important to understand loading characteristics to prevent potential voltage and thermal excursions in the short term, and to inform network planning [2] in the long term. Adopting a monitoring regime similar to that used at the transmission level i.e., recording and storing historical values, is prohibitive owing to the number, size and heterogeneity of the LV networks - tens of thousands of 11 kV substations in a typical GB DNO licence area [3], and in an urban conurbation such as London, there are 19,583 LV substations [4]. Even if costs for sensing and communications infrastructure for each substation or network bus installation are fixed, consideration must be paid to the OPEX costs, such as data transmission and storage costs.
Scenario generation is one viable route to forestall potential problems from the introduction of new technology or general load growth. However, a baseline performance measure is difficult to obtain due to the lack of LV observability. At the distribution level, the primary concern is the extremes: headroom for thermal exceedances and footroom for voltage violations [5]. Typical Load Profiles (TLP) exist [6], but these are generally seasonal, not weather sensitive and lack the diversity that would enable a cohort of premises to be aggregated together into a plausible LV substation load. In defining plausibility, the peak demand sharpness (and under embedded PV scenarios, trough depth) should reflect the range of time use across neighbourhood premises: too sharp would result from near identical behaviours and hence an overestimate of potential threats; too blunt assumes greater diversity than is likely, and therefore a potential overestimate of headroom. Generative modelling from a limited set of historical exemplars offers a potential solution, but the varied nature of LV substation loading introduces its own set of complexities and hurdles. The statistical distribution of LV substation load does not follow any common probability density function [7], making sampling approaches unrepresentative unless carefully chosen for particular locations. Markov Chain approaches also struggle with capturing aggregated time use resulting from routine behaviours [8]. Understanding the present loading characteristics is essential as this condition could then further worsen with the extensive installation of distributed generators as well as low-carbon heat and transport [9]. This will cause equipment to operate under atypical conditions that impact the wider network [10], potentially up to the transmission level. For a Distribution Network Operator (DNO), this insight is necessary for long-term planning of reinforcements that will facilitate LCT adoption and generally increase resilience over time periods of months, years and decades. For a Distribution System Operator (DSO), the same insights will yield short-term generation dispatch and load curtailment potential within the constraints of the physical network infrastructure, with both active and reactive turn-up and turn-down capabilities articulated accurately via a flexibility envelope  [11]. This, in turn, will articulate substation flexibilities that can provide balancing at the transmission interface, which, combined, represent regional-scale scenarios from fundamental end-use insights - unlocking this knowledge of aggregating demand would enable better system planning as well as policy design over a longer period.
LV distribution network loads are much closer to individual premises end use, meaning that lower levels of aggregation will reflect highly variable behaviour routines which may result from social norms, variation in appliance ownership and specification, and also reactions to localised weather conditions. Previously, LV substation data was modelled using conventional methods such as aggregated average profiling [6] with a diversity correction. While this approach was historically sufficient, the increasing penetration of LCTs now necessitates more advanced and accurate methods. To approach this complex and non-stationary behaviour, it is proposed here that a generative AI model is applied, which is capable of synthesising both Active and Reactive power data for LV substation load data based on a set of observed influencing factors, namely, weather and calendar variables. This is novel in the following ways: Firstly, applying a deep generative model to LV substation load modelling. More specifically, a Diffusion model, which is the state-of-the-art method for the generation of time series data. Deep generative models can learn underlying temporal and statistical correlations within the data and make use of substation metadata along with weather and calendar variables, enabling higher-quality load profile scenario generation. Secondly, this study considers the synthesis of reactive power profiles at the LV substation level. Reactive power is a key part of load behaviour and technologies such as heat pumps; the assumption of a flat power factor is not realistic, as specific appliances generate a reactive power requirement, the use of which is driven by behavioural routine. Coupling of routine behaviours means that this relation is not necessarily 1-to-1. Furthermore, the synthesis of reactive power facilitates a power systems analysis. Unlike in other similar works, the evaluation of realism is not entirely down to statistical measures – co-occurrence of load behaviours across cohorts of synthesised LV substation will affect the load and voltage profile at higher levels of the network, hence a power system evaluation needs to be carried out to ensure that diversity (and synchronisation of premises or substation extreme events) is realistic. The proposed model can capture and replicate realistic substation-specific behaviour without the significant overhead of mass data collection and storage.
This paper is structured as follows: Section 2 reviews the related work on load synthesis tasks. Section 3 proposes a generative approach to LV substation load data synthesis based on contextual conditions. Section 4 evaluates the performance of the model through multiple synthesis metrics and visualisations. Section 5 conducts a Load Flow analysis to evaluate wider power system coherence through the substation hierarchy. Section 6 presents the conclusions and future work of the study.
2 Related Work
Approaching an energy system in transition, it is important to be able to identify the capabilities and limitations of infrastructure already in place, with its present-day utilisation. Understanding scenarios in the present allows the ability to accommodate new technologies to be anticipated. In the past, methods of average profiling have been used to model the load behaviour of electricity end use, such as system operator typical load profiles (e.g  [6]) or clustering approaches [12, 13]. These methods leverage highly generalised ‘typical’ customer behaviours to provide a series of averaged profiles which incorporate factors such as building type, seasonal trends, days of the week, etc. However, these methods remain static and lack diversity when modelling the collective load on the substation in which a cohort of them is connected. Realistic modelling requires an approach that captures the diversity inherent in behaviour profiles as opposed to those generated via coarse-grained averaging.
One approach is Gaussian Mixture Models (GMMs) [14, 15, 16], that involves fitting a mixture of Gaussian distributions to model load profiles based on a number of components (the number of Gaussian distributions fitted across the dataset). The optimal number of components can be calculated using the model complexity metric, Bayesian Information Criterion (BIC), to provide the best fit for the given dataset. Although GMMs improve upon the traditional average load profile approaches by capturing more realistic load profiles, they lack the temporal and statistical accuracy essential for LV substation load monitoring.
Deep learning approaches to load synthesis have been proposed for smart meter data through the use of Generative Adversarial Network (GAN) models and Convolutional Variational Autoencoders (CVAEs) [17, 18, 19]; there are variants whereby these deep learning models have been combined with clustering methods [20, 21]. More recently, applications of Diffusion models have been applied to power systems data; however, not in the context of load modelling at the LV substation level. Source–load scenarios for renewable generation are synthesised at a power systems level using diffusion models [22, 23]. Diffusion models have, however, been applied to load modelling at the smart-meter level of the network [24], and net-load synthesis has also been applied to customer data [25].
Despite the similarities between smart meters and LV networks, these approaches are not directly applicable to LV networks due to the broader range of factors that must be considered, such as weather and LCT penetrations, for example. These considerations result in load profiles at different levels of the network having different underlying statistical distributions. Furthermore, although the studies have shown excellent ability to synthesise individual smart meter load profiles [17, 18, 19, 20, 21], they have no consideration of diversity across cohorts of premise meters. Consequently, the quality of the aggregated profile at the next voltage level up may lack representativeness. Additionally, these models are often developed to serve different objectives such as increasing data quantity due to availability issues stemming from privacy [17, 18, 19, 20, 21]. The purpose of the proposed study is to synthesise load profiles based on individual characteristics rather than just typical scenarios. The addition of extra conditions (e.g., daily min, mean and max power, and substation-specific metadata) alongside the typical conditions observed in the literature (weather and calendar variables) allows the new model to fulfil this specific goal  [17, 18, 21].
The proposed conditional diffusion model permits load profile synthesis at low aggregations on distribution networks with conditions such as weather, statistics and calendar variables that address the challenge posed by diversity when utilising average profiling approaches. Furthermore, the ability to synthesise reactive power enables load flow analysis, which allows evaluation of wider power system coherence.
3 Methodology
This section proposes the diffusion models developed to synthesise load profiles with an aim of providing operational data with realistic variability dependent on various factors such as weather data, calendar variables, and behavioural trends. Firstly, the concept of diffusion models and the SSSDS4 model is introduced in Section 3.1, then the developed models are presented in Section 3.2, and the dataset used is described in Section 3.3. Finally, the evaluation metrics and visualisations used are outlined in Section 3.4.
3.1 SSSDS4 Diffusion Model
Diffusion models were chosen as the generative approach for synthesis in this study owing to their excellent performance in imputation and generation across various data modalities. Furthermore, diffusion models have previously outperformed GANs and CVAEs specifically in time series data [26, 27, 28, 29, 30].
Figure 1: Diagram outlining the forward and reverse process of the diffusion model, key variable names are provided with the diagrams.
Diffusion probabilistic models generate new independent samples that statistically resemble the distribution of the underlying data on which they are trained. They operate in a forward (training) and reverse inference process. The diffusion forward process follows a Markov chain where Gaussian noise is progressively added to a sample $x_{0}$ in $T$ steps; $t=1,\ldots,T$. The level of noise $z(t)$ added in each step $t$ is typically drawn with a linear noise scheduler $\alpha$ between an upper and lower noise level $\beta_{0}\,\text{and}\,\beta_{t}$ resulting in a corrupted sample $x(t)$. During training, rather than simulating the multiple steps of the Markov chain for every training iteration, it is more efficient to corrupt the sample in a single stage for all steps $t$ by sampling directly from cumulative Gaussian noise, which can be calculated using $\alpha$ [30]. Figure 1 illustrates the forward process of the model, where for a randomly sampled diffusion step $t\in(0,T]$. The model corrupts the sample with $z(t)$ and predicts $\epsilon_{\theta}(x_{t})$, the level of the cumulative noise in $x_{t}$ which accounts for the noise added in all steps $0\to t$. The model is trained for multiple iterations and learns to predict noise for all values of $t\in(0,T]$. During inference, the reverse process recovers the sample $x_{0}$ from $x_{T}$, where the latter is pure Gaussian noise. The inference process also follows a Markov chain but in the reverse direction, from steps $t=T,\ldots,0$. The noise is iteratively removed by first predicting the cumulative noise from the current step $x_{t}$ to $x_{0}$ - to the beginning of the chain $\epsilon_{\theta}(x_{t})$. Then the model removes the portion of the noise attributed to step $x_{t}$ to $x_{t-1}$, which can be estimated using Equation 1. Finally, at each step, a variance term is added to $x_{t}$ to retain stochasticity in the generation process and permit synthesis of diverse samples [31, 32].
$x_{t-1}=\frac{x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\cdot\varepsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}}+\sigma_{t}$ (1)
The proposed diffusion model is built upon an implementation of the SSSDS4 model [26]. SSSDS4 provides a time series implementation of the diffusion model enhanced with the combination of structured state space models, to allow capturing long-term dependencies in the data, such as seasonal and diurnal trends linked by weather and behavioural routine [33]. The SSSDS4 implementation of diffusion models was chosen due to its state-of-the-art performance in several generation tasks across multiple datasets when benchmarked against other diffusion-based approaches [26]. The SSSDS4 model was adapted to utilise conditional inputs to constrain the generation process and synthesise samples from specific areas of the overall distribution. Conditions are introduced to the model as additional channel inputs and utilise the binary masking mechanism to discern signal and conditional inputs. When the mask value is 1, the original input is provided to the model as a condition; when it is 0, the input is replaced with noise, and the model must learn to denoise it. The model then learns a transformation from the conditional inputs and the signal noise vector to a novel load profile.
Furthermore, the SSSDS4 model source code contained an implementation error in the training loop relating to the handling of conditional inputs. Algorithm 1 outlines the modified training step in the SSSDS4 model. There is one key difference between this and the previous model. Line 7 restores the conditional values after the entire sample is corrupted with noise. Previously, this was handled differently; the conditional values were substituted into $x_{std}$ (the noise vector). Then, when the sample is corrupted, the conditional values would not be corrupted with noise. However, the noise corruption equation outlined in line 6 only balances the data for values of $x_{0}$ $\in[0,1]$. In many cases, this is not sufficient; a more effective method is to corrupt the entire sample with noise, then replace the values after corruption. This aligns with the behaviour during inference, where the conditional values are replaced at each step.
Algorithm 1 SSSDS4 Modified Training Step
[…]
12:update_parameters(loss)
13:
Figure 2 shows a diagram for the implementation of the SSSDS4, highlighting the model inputs/outputs and key layers. The minor modification made to the model can be observed in the inputs and conditions being passed to the model. Figure 3 provides more detail of the residual block section of the diagram. The block represents the structure of a single residual layer, and multiple of these blocks are chained together. After a hyperparameter search, 36 residual blocks were used. The output of all residual blocks is then added together and fed to the subsequent 1D Convolutional layer.
Figure 2: SSSDS4 Model implementation outlining model inputs, outputs and key layers/blocks.
Figure 3: Diagram of the Residual Block within the SSSDS4 Model.
3.2 Load Profile Synthesis Models
This study proposes three diffusion models:
  1. 1.
    LVGenU - The unconditional diffusion approach provides a blind synthesis without utilising conditions. In this form, the diffusion model attempts to recreate load profiles from Gaussian noise alone.
  2. 2.
    LVGenWC - This model adds conditional variables which require no daily storage; thus, more accurate daily loads could be reconstructed without the requirement of any recorded values. The features used for this model are as follows; calendar information (day of week, month etc.) which can be directly inferred from the timestamp, weather forecast (temperature, humidity, wind speed, etc.) collected from API based on the substations latitude/longitude co-ordinates, and substation information (number of customers connected) which will be stored for each substation separately.
  3. 3.
    LVGenWCS - The next addition of cues involved passing the daily minimum, mean, and maximum of the active and reactive power. This allows for a significant improvement of synthesis and reconstruction, particularly in the more extreme values from substations with unique behaviours.
Each model was trained with the optimal neural network configuration defined in [26], with 200 diffusion steps. Each model was trained with a maximum of 200 epochs, but was halted earlier if test loss plateaued. Figure 4 shows the convergence of each model based on the MSE vs the number of epochs; LVGenU converges relatively quickly because there are no conditions for the model to learn from; LVGenWC and LVGenWCS require a significantly higher number of iterations before the models converge.
Figure 4: Test Loss for each Model through epochs.
3.3 Data Set Description
Distribution substation monitoring data was acquired from the National Grid Electricity Distribution (NGED) OpenData [34] platform. It constitutes a subset of monitored LV Networks consisting of 1,431 substations, which total to approximately 19,050 days of data (as of 6th of June 2024). Measurements include the active and reactive power for each phase (L1, L2, L3). Data for each substation was also paired with its associated metadata (such as primary substation number, number of connected customers, LCT capability, etc.), weather from the Meteostat API [35] and calendar information. Where there were missing or incomplete data samples, a discard strategy was adopted. Data cleansing also removed outliers such as data with unrealistically high values (greater than 1,000 kW), incomplete days, and other anomalous data (such as constant values recorded throughout the day), resulting in a cleansed dataset of 13,756 days. The dataset was split into training and testing sets based on the number of observations per substation, placing substations with a low number of observations in the testing set. This was to inhibit overfitting due to limited training data. A cut-off of 10 days was chosen to achieve a balanced split, with observations having fewer than 10 days placed in the testing set. This threshold was selected to maintain an approximate split of 70% for training and 30% for testing. Furthermore, this method of splitting was also chosen to ensure representative data synthesis at the substation level; data segments from the same substation could not be included in both the training and testing sets, i.e. data leakage was prevented by allocating all data from a particular substation to either the test or train sets.
The number of customers served by a substation was captured through fixed-width bins: 0-99, 100 - 199, etc.; the final bin contained all substations with greater than 600 customers. This form of discretisation was performed to prevent the model from memorising specific behaviours of individual substations based on their associated number of customers. This, in turn, enables the model to generalise and scale to potentially larger network scenarios. To create representative training/test sets, the split was performed in a stratified fashion based on the discrete group of the number of customers. Finally, in order to permit testing of the load flow simulation presented in Section 5, a set of substations which are connected to the same Primary Substation were placed in the testing set.
3.4 Evaluation Metrics and Visualisations
In this study, two metrics are used to evaluate the performance of the data synthesis models in two key areas. Mean Squared Error (MSE) is used to measure the reconstruction errors of the generated load profiles vs the actual. MSE was chosen as it is an error metric that is routinely used with machine learning applications. The Maximum Mean Discrepancy (MMD) is used to measure the quality of the synthesis. MMD is a distance-based metric measuring the difference between real and generated probability distributions by comparing their means in a high-dimensional feature space [36]. A lower score represents a more similar data distribution. MMD was chosen as it is regularly used in data synthesis studies as a distance-based metric [37]. Additional metrics, inspired by previous research [38, 39], were also considered and are listed in the table. These further distance-based metrics Wasserstein distance [40], marginal score and MiVo [38]. The remainder of this section focuses on MSE and MMD metrics, with the additional metrics referenced only when they provide additional insight.
Data visualisations are used alongside these metrics to further evaluate the quality of the generated samples. The first kind of visualisation used is distribution plots, which contrast the implied probability distribution of the actual data and each proposed model. A zoomed-in section of the tail of the distribution is also provided to inspect the important extreme values, which in operational scenarios could result in problems such as voltage collapses. The second method of visualisation is demonstrated through decile plots, which plot the values of the load at various deciles of the implied probability distribution. The final visualisations used are Autocorrelation Function (ACF) plots, which measure the correlation of the data with a lagged version of itself across various time intervals. Similar ACF plots indicate that datasets share a similar underlying temporal structure, meaning they exhibit comparable diurnal patterns or intra-day dependencies. Thus, the resemblance in ACF plots implies that the datasets may have been generated by similar processes or are subject to similar temporal influences such as behavioural routine or local weather conditions [41, 19].
4 Synthesis Results
This section describes the performance of the proposed models through the use of various metrics and visualisations. Proposed model results are benchmarked against a Gaussian Mixture Model (GMM) and a Wasserstein GAN (WGAN). These benchmarks are mainly used for comparison with the unconditional diffusion model, after which conditional diffusion models are introduced to demonstrate the enhancements achieved through conditioning. In the subsequent section, the Tao Vanilla model is added as a further benchmark. This model could not be applied in this section as it must be trained on the specific target substations. Given that the test set consists of unseen substations, it is not possible to train the Tao vanilla and provide a fair comparison.
Random samples from GMMs provide a good model to compare metrics against the diffusion approach due to their generative nature and their previous use in load synthesis tasks [14, 15]. GMM can be fitted against the data set distribution using the optimal number of components derived from the Bayesian Information Criterion (BIC), which for this data was one component for the active power and four for the reactive power. The samples can then be generated by drawing samples from the resulting GMM at random.
A Wasserstein GAN (WGAN) is trained as a comparable deep learning approach. Information regarding the training methodology and procedure is described in A.
Metric GMM WGAN LVGenU LVGenWC LVGenWCS
MSE 0.93 1.3 0.68 0.29 0.1
MMD 0.54 0.67 0.44 0.13 0.005
Marginal Score 0.24 0.09 0.16 0.12 0.06
MiVo 0.73 0.97 0.57 0.21 0.05
Wasserstein Distance 0.29 0.28 0.37 0.15 0.015
Table 1: Error metrics for each model. Results are calculated on the entire unseen test set using scaled values. Best scores are highlighted in bold.
Table 1 calculates the score of each metric for each model. When comparing unconditional models, LVGenU outperforms the GMM and WGAN in synthesis and reconstruction with lower MSE and MMD scores. GMM and WGAN do exhibit a lower Wasserstein distance, with WGAN further demonstrating a significantly lower marginal score. The LVGenU is trained on MSE and consequently does not pay direct attention to the distribution shape. On the contrary, the WGAN contains the Wasserstein distance in its loss function and hence the model is directly optimised to capture the distribution shape without considering the temporal coherence of the samples (as it will be discussed later when examining Autocorrelation Function performance; Figure 7). Generative models that utilise conditions in LVGenWC and LVGenWCS significantly improve all these scores, with an extremely good match between the synthesised and real profiles.
(a) Active Power
(b) Reactive Power
Figure 5: Comparison of the distributions of Active and Reactive Power for different models and the real data. The detail shows how the upper tail behaviour is captured.
Figure 5 plots the active and reactive power distribution curves of the real dataset and each of the proposed models. Results show a similar trend to what is observed in the Marginal score and Wasserstein Distance; the increase in conditions results in a more accurate capture of the distribution. The GMM provides a good capture of the overall distribution for the active and reactive power; however, it fails to capture both tails of the distribution. This is due to the GMM being unable to capture the higher-order moments. The WGAN demonstrates a better overall fit to the distributions compared to the GMM, mainly due to its improved capture of the peak of each distribution. However, it still fails to accurately capture the lower tail. This behaviour reflects the influence of the WGAN’s penalty term, which contributes to its improved Wasserstein Distance and Marginal Score. However, this emphasis results in weaker performance in all other metrics, namely MSE and MMD. LVGenU demonstrates a pessimistic generation rarely diverging from the median samples in order to minimise the MSE loss function. The lower MSE allows the model to achieve better reconstruction of load profiles; however, this comes at the cost of a less accurate distribution capture, highlighted by a higher Marginal score and Wasserstein distance compared to the GMM and WGAN. LVGenWC improves upon LVGenU by generating a wider variety of samples and achieving a more accurate capture of the distribution, as demonstrated by an improvement in each of the distance-based metrics. This is due to the distribution containing a less significant oversample of the median, and not as severe a miss of the extreme values. However, both issues persist, though not to the same extent as with LVGenU. The distribution generated by LVGenWCS (orange) almost perfectly matches the real distribution (red), making it difficult to distinguish between the two in the figures, as LVGenWCS closely overlaps the real distribution. The model is also able to capture multi-modal behaviour in the tail of the distribution for the reactive power. This leads to a substantially lower MMD score compared to all other models, representing the most significant improvement in the metric by a wide margin.
Figure 6: Decile plots for the 10th, 50th, and 90th decile for each point of the daily load. Plots show the real data, and each model’s generated samples for active and reactive power. Results show how the general trend of the different distributions is captured by each model.
Figure 6 contains the active and reactive power decile plots for the real dataset and each of the proposed models. The figure plots the typical load profiles throughout various deciles for active and reactive power. Results show the GMM misses the real load profile at every decile for active and reactive power, all to a similar degree; these consistent misses result in a high MSE. WGAN is in the correct range for the 90th decile, but misses the shape of the load profile. In the other deciles, WGAN misses more severely than the GMM, resulting in an even higher MSE due to these extreme missed values. LVGenU is much closer to the real data for active and reactive power at the 10th decile compared to GMM and WGAN. At the 50th decile, similar results to the GMM are obtained. The model completely misses the 90th decile, however still manages to retain a lower MSE when compared to the benchmark models. The generally high MSE for each unconditional model results from their uninformed generation process, as neither model has any indication of the load profile it is attempting to reconstruct (guidance that subsequent models receive through their conditions). LVGenWC provides a more accurate capture of active and reactive power at the 10th and 50th deciles, performing comparably to LVGenWCS in these areas. However, it struggles to capture the 90th decile, a trend similar to what was observed in Figure 5. This is reflected in the MSE, where LVGenWC shows improvement due to its better capture of the 10th and 50th deciles, along with a less severe miss at the 90th decile. LVGenWCS further improves the MSE by achieving a much better capture of the 90th decile for both active and reactive power.
Figure 7 contains the active and reactive power plots for the ACF. Results show all diffusion models are very closely correlated with the real data throughout the lag values for the active power. The GMM loses correlation with the real data for active from lag 25 onwards. For reactive power, the WGAN in particular struggles and loses temporal coherence with the real data. For the other models, including the GMM, the correlation is strong. The MMD metric can provide additional insight into the similarity between the temporal structure of the real and generated data sets. GMM has a lower MMD than WGAN due to WGAN’s poor performance on the reactive power. LVGenU further improves upon the MMD performance of GMM, which may be attributed to improvements observed in the ACF plot for active power. The addition of conditional inputs to the diffusion model further improves the MMD, similarly to all other metrics. This highlights the importance of the conditional values for the generation process.
(a) Active Power
(b) Reactive Power
Figure 7: Comparison of ACF plots for active and reactive power data.
Together, the metrics and visualisations outline the performance of each model and demonstrate the improvement from the baseline GMM and WGAN to a diffusion model in an unconditional perspective. This is through the ability of the diffusion model to capture higher-order moments in the data and better replicate the temporal structure of real data. Then the addition of conditional inputs led to further improvements being observed in the results. The LVGenWC model shows promise as a scenario generation tool, but does not adequately sample from the extremes of the distribution, nor does it consistently recreate load profiles to high accuracy. To address this limitation, additional inputs or modifications to the model may be required to capture these extreme values effectively. However, for the specific goal of this study, LVGenWCS is far more suitable as it can generate specific load profiles from individual substations to a high quality and accuracy right across the distribution. In subsequent sections, it is demonstrated how the models generated aggregated load profiles perform in a wider power systems context through the application of power systems models. Although informative in terms of model performance, synthesis metrics alone are not sufficient. A model must also produce robust phase angles and voltage estimates across the entire network to serve as a credible replacement for rigorous monitoring.
5 Power System Analysis Case Study: Urban Scale Load Flow
Results from the Diffusion models, when applied to the synthesis of LV loads, show promising accuracy; to evaluate if generative data could replace real load data, analysis of single loads is insufficient for assessing overall network stability and safety. One important consideration is to ensure that the voltages are within statutory limits to prevent any over or undervoltage conditions. Furthermore, it is important to determine the extent to which the network components, like transformers, cables, or capacitors, may need to be reinforced or upgraded based on the load flow analysis.
To address this, simulations of Medium Voltage (MV) distribution feeders are performed, which are based on representative urban and rural 33kV networks in Great Britain (GB). The feeders are populated with actual LV monitoring data, and load flows are calculated to obtain bus voltage magnitudes and phase angles. Additionally, the ‘Tao Vanilla’ model [42] is provided as an additional benchmark model alongside the GMM and WGAN. The Tao Vanilla model is widely used as a benchmark in power system case studies and aligned research to conduct load forecasting and analysis of load flow. The model uses multiple linear regression to predict the load based on a set of load-driving instantaneous criteria, including weather and calendar data:
${\rm E}(Load)=\beta_{0}+\beta_{1}\times Trend+\beta_{2}\times Day$ (2)
$\times Hour+\beta_{3}\times Month+\beta_{4}\times Month$
$\times TMP+\beta_{5}\times Month\times TMP^{2}+\beta_{6}$
$\times Month\times TMP^{3}+\beta_{7}\times Hour\times TMP$
$+\beta_{8}\times Hour\times TMP^{2}+\beta_{9}\times Hour$
$\times TMP^{3}$
Tao Vanilla can express the relationship between load, temperature, and seasonal and diurnal covariates through multilayer linear relationships, making it a good benchmark model for expressing these simple relationships.
5.1 United Kingdom Generic Distribution System (UKGDS)
UKGDS contains several models that can represent the behaviour of the GB distribution networks. In this paper, a UKGDS 77-bus test network is examined. It is simulated for an urban area with a high customer density, containing both 33 kV and 11 kV networks. The structure of the UKGDS network is shown in Figure 8.
Figure 8: UKGDS Network Test Network used to assess MV impact from LV substation load behaviour.
The network covers an area of approximately $10\,\text{km}^{2}$. It consists of one 33 kV substation that acts as the slack bus, with two transformers at the same location, stepping down the voltage from 33 kV to 11 kV. In total, there are 75 loads connected to the remaining 75 11 kV substations distributed throughout the network. The average distance between each bus is about 0.75 km. The network connections are organised into several levels, which enable distant substations to operate at lower voltages. In this network, the designed limit for the voltage magnitude is between 0.97 to 1.03 p.u., and the base capacity is 10 MVA. In this study, the 75 LV substations were analysed using three repeated sets of data consisting of 26 actual LV substations to simulate network performance. This approach aims to maximise the diversity of load flow results across the network. Additionally, only the 26 substations that were active in the network were also used for testing. The results demonstrated similar results to those of the 75 substations, but showed reduced diversity on the low-voltage side, attributed to the lower load within the network.
5.2 Load Flow Setup and Analysis
The load flow analysis is used to ensure that the power system operates within its network limits of 0.05 p.u. for voltage magnitude and 10 degrees for phase angle, which are standard values in GB distribution networks. Thus, determining the most economical generation dispatch [43]. The load flow used is based on the Newton-Raphson method and is performed using the Python library pandapower [44]. Given an $n$-bus system with one slack bus, which has a constant 1.0 p.u. voltage magnitude and zero phase angle. The complex power at the node $n$ can be calculated as [45]:
$(P_{n}+jQ_{n})=\overline{E}_{n}\sum_{m=1}^{n}\overline{Y}_{nm}\overline{E}_{m}$ (3)
where $P$ and $Q$ are the active and reactive power injecting into the bus, $E_{n}$ is the node-to-datum voltage, $Y_{n}$ is the element of the admittance matrix, j is a complex number, and $\overline{E}$ indicates complex quantities. To apply the Newton method, a Jacobian Matrix is used to represent the partial derivatives of the load flow equations, which can be represented as:
$J=\begin{bmatrix}\frac{\partial P}{\partial\theta}&\frac{\partial Q}{\partial\theta}\\ \frac{\partial P}{\partial V}&\frac{\partial Q}{\partial V}\end{bmatrix}$ (4)
where $V$ is the voltage magnitude and $\theta$ is the phase angle, the voltage magnitude and phase angle can be iteratively updated using the Jacobian matrix by:
$\begin{bmatrix}\Delta\theta\\ \Delta V\end{bmatrix}=-J^{-1}\begin{bmatrix}\Delta P\\ \Delta Q\end{bmatrix}$ (5)
Therefore, the updated voltage magnitude and phase angle after each iteration is:
$\theta^{(k+1)}$ $=\theta^{(k)}+\Delta\theta$ (6)
$V^{(k+1)}$ $=V^{(k)}+\Delta V$ (7)
The iterations continue until the power mismatches and ($\Delta P$ and $\Delta Q$) are within a specified tolerance. In this analysis, only the voltage magnitude and phase angle are considered, as these are the main parameters of interest to the DNO. Although line losses, thermal margins, and line currents can also be included in power flow studies, the focus here is on assessing whether the synthesised load can represent the network operating conditions. Since the DNO’s main concern is whether these effects are properly captured by the power flow analysis, it is well established that the voltage magnitude and phase angle can be presented as key indicators of network performance [46]. Other aspects can be explored in future research by generalising the contribution.
5.3 Load Flow Simulation Results
Figure 9: Comparison scatterplot of Voltage Magnitude predictions versus ground truth for each model.
Table 2: Comparison error table of Voltage Magnitude predictions versus ground truth for each model.
Model MAE (V) $R^{2}$ 5th Percentile Error (V) 95th Percentile Error (V)
Tao Vanilla 3.13 -2.48 0.536 6.8
GMM 3.31 -0.37 0.184 9.0
WGAN 5.28 -2.96 0.633 13.5
LVGenU 7.03 -5.47 2.605 15.4
LVGenWC 4.83 -2.57 1.606 10.6
LVGenWCS 0.75 0.89 0.030 2.720
Figure 9 and Table 2 present a comparison of the voltage magnitude results obtained from the benchmark and proposed diffusion models. Models tend to perform well in estimating higher voltage magnitudes when the system load is at a minimum, but tend to be less accurate at lower voltage magnitudes when the load is near its peak. In general, most methods overestimate the voltage magnitude, which could result in inappropriate decisions by system operators. In contrast, the LVGenWCS method consistently captures the trend of the voltage magnitude accurately, regardless of whether the values are high or low.
The results in Table 2 compare the six methods, which show that LVGenWCS demonstrates a MAE that is five times lower (than the next best model). In addition to MAE, the 5th and 95th percentile errors are also reported. The range between the 5th and 95th percentiles provides insight into the distribution of the majority of errors, effectively excluding the most extreme 5% at either end. Such insights are especially valuable for power system operators, whose primary concern is the reliable and consistent performance of predictive models under typical operating conditions. The LVGenWCS method also exhibits substantially lower errors within the 5th to 95th percentile range, indicating greater accuracy and reliability across the full spectrum of voltage magnitudes. Moreover, the results suggest that, in the context of a constrained power system where the system operator aims to maintain the voltage magnitude within ±0.05 p.u., only the LVGenWCS method possesses the capability to provide load predictions that remain within this limit. Consequently, it is the only method that can reliably support the system operator in making appropriate operational decisions.
Figure 10: Single LV bus time series of hourly average voltage magnitude demonstrating temporal fidelity between real and LVGenWCS data with min and max thresholds.
Furthermore, Figure 10 compares the LVGenWCS with the real data for a single bus test case. The LVGenWCS model was run ten times, and the results were averaged to produce a representative output, with confidence intervals also reported based on the model’s upper and lower outputs at each timestep. The findings demonstrate that the LVGenWCS model provides reliable results that replicate the dynamic behaviour of the network. The LVGenWCS model closely aligns with the real load data, accurately capturing voltage patterns with reduced deviations in amplitude. Although some of the troughs are overestimated in certain model results, the average of the results is sufficiently accurate, with an acceptable error margin. This indicates that the LVGenWCS model offers reliable accuracy for the network and serves as a good fit for predicting voltage magnitudes.
Figure 11: Comparison scatterplot of Phase Angle predictions versus ground truth for each model.
Table 3: Comparison error table of Phase Angle predictions versus ground truth for each model.
Model MAE (deg) $R^{2}$ 5th Percentile Error (deg) 95th Percentile Error (deg)
Tao (Vanilla) 0.056 -2.80 0.02 0.09
GMM 0.028 0.19 0.005 0.06
WGAN 0.041 -0.89 0.007 0.08
LVGenU 0.046 -1.25 0.023 0.08
LVGenWC 0.029 0.01 0.01 0.05
LVGenWCS 0.007 0.93 0.0004 0.02
Figure 11 and Table 3 present a comparison of the Phase Angle results obtained from the benchmark and proposed diffusion models. The results show the Tao Vanilla model performs poorly as it is not designed to predict the reactive power. The results follow similar trends to those observed with voltage magnitude, where models excluding LVGenWCS provide acceptable results for higher values, but display significant bias errors for smaller values. The LVGenWCS model is consistently able to capture both higher and lower values, although the error in smaller values is greater than what was observed with the voltage magnitude results. This indicates that reactive power is considerably more difficult to predict.
The results in Table 3 also show that the LVGenWCS method outperforms the other methods, again demonstrating MAE that is roughly five times lower than the next best model, and ten times lower for the 5th percentile. The reactive power values here are considerably low due to the data only consisting of resident loads. The results indicate that only the LVGenWCS method shows potential for wider application in reactive power prediction for networks containing commercial or industrial loads, thereby providing system operators with more reliable day-ahead predictions.
Figure 12 compares the phase angles over time for a single bus between real values and the LVGenWCS model. The results again show similar trends to the voltage magnitude analysis. The LVGenWCS model results closely follow the trend of the real data, effectively capturing both amplitude and frequency of the oscillations. The LVGenWCS model exhibits good deviations from the real data, particularly at the peaks and troughs.
Figure 12: Single LV bus time series of hourly average phase angle demonstrating temporal fidelity between real and LVGenWCS data with min and max thresholds.
Overall, the strong correspondence between the real data and LVGenWCS model outputs for both voltage magnitude and phase angle highlights the effectiveness of the LVGenWCS model in replicating complex voltage characteristics based on the load profile. This high level of accuracy suggests that the model can be reliably used for both prediction and analysis in distribution power systems.
6 Conclusion
Transitioning towards Distribution System Operation in legacy power systems requires understanding of load interactions at the sub-regional level to avoid thermal and voltage constraint violations. Extensive monitoring would ultimately inform mitigating strategies to a high standard, but it is prohibitively expensive to deploy and maintain comprehensively across distribution networks owing to the sheer volume of assets. While load behaviour synthesis offers a convenient alternative, demand characteristics at low voltage substations must be realistically diverse so that higher voltage level behaviour is accurately represented: lack of diversity results in voltage collapses and thermal exceedances, excessive diversity will excessively smooth out any extremes. To address the problem of realistic distribution load profile synthesis, this work has contributed a Generative Conditional Diffusion Model for reconstructing LV substation load profiles based on a minimum number of cues. Realism was demonstrated on an individual level through the validation of temporal and statistical characteristics, while coherence in the wider power system context was demonstrated through the propagation of synthesised LV loads through a representative MV network. The load flow results demonstrate that the synthesised LV substation profiles have a similar impact on the network when compared with results obtained from using meter data. The implication of the equivalent load flow result is that diversity across LV substations synthesised by the diffusion model has temporal variability, which reflects load diversity appropriately in most cases. This model provides a good synthesis of the substation base load profiles, which is necessary for power system studies and is a prerequisite to benchmarking the impacts of LCT penetration. Future development will entail retraining the model with reliable LCT metadata, such as the proportions of the distributed energy sources adopted under the LV substation, such as photovoltaic (PV), wind and EV charging, to better learn and replicate the effects of these diversity changing loads. From a behavioural perspective, understanding rebound effects from time of use tariffs (for instance, extrapolating [20]), there is also the potential for model trust to be questioned, which motivates the need for future work in model explainability to account for the plausibility of what has been captured by the model and how it produces subsequent load representations on which planning and operational decisions are made.
Acknowledgments
Results were obtained using ARCHIE-WeSt High Performance Computer (www.archie-west.ac.uk).
Appendix A WGAN Training
This section will cover the training of the Wasserstein GAN (WGAN) model used as a benchmark in the study. The WGAN model used was implemented by Hao et al. [47]. Like with SSSDS4, the model was used with existing hyperparameters with one minor modification. A non-autoregressive generator was added to focus the model on synthesis rather than forecasting.
Generally, training of a GAN can be challenging to know when to stop the model based on its multi-objective functions  [48]. Specifically in the WGAN, where the extra loss penalty can act as another optimiser. For training, the model was tested to 100 & 200 epochs and also stopped when visual inspection of the loss terms looked favourable. The deviations in results between each other were minimal, but the earlier-stopped model provided slightly improved results. Figures 14 & 13 show the loss terms during training. Given the unconditional nature of the model the training converged quite quickly once the Wasserstein penalty had stabilised; there was little extra information for the model to learn. Out of the scenarios described above, the best results were obtained when stopping the model around 2000 iterations, when the generator’s performance was better. However, it must be noted that the difference in final results and metrics is minimal.
Figure 13: Loss terms for the Generator and Critic for each training step.
Figure 14: Wasserstein Loss penalty term for each training step.